OpenStudy (anonymous):

Samuel works for a high-speed rail company that wants to develop a new rail line. Samuel’s project is to find a train that is the second fastest in the world. The Shinkansen Bullet Train in Japan is reported to go as fast as 320 kilometers per hour. The TGV train in France can reach speeds of 89.44 meters per second. Explain to Samuel how to find what speed his new train must go to be the second fastest in this group. Then, find an appropriate speed for Samuel’s train in miles per hour and use the formula distance = speed • time to find how far the train can go in 2.5 hours. Use complete sente

OpenStudy (anonymous):

I did this problem ur flvs right

OpenStudy (anonymous):

?

OpenStudy (anonymous):

flvs ?

OpenStudy (anonymous):

I'll try and help u

OpenStudy (anonymous):

ty

OpenStudy (anonymous):

lol u changed the name from ellen to samuel

OpenStudy (anonymous):

im doing a online summer school

OpenStudy (anonymous):

well its simple...

OpenStudy (anonymous):

convert 320 kilometers per hour to meters per second

OpenStudy (anonymous):

urmmm

OpenStudy (anonymous):

Yes mark flvs right its online im doing algebra 1 in the ummer goin into 8th

OpenStudy (anonymous):

here i did this last summer ill help u hold on

OpenStudy (anonymous):

austin the school is in gresham

OpenStudy (anonymous):

First you have to convert all the speed in the same unit.. its entirely upto you to decide.. let us convert speed of SBT train in Japan to m/sec

OpenStudy (anonymous):

now to convert it we multiply it's speed by 1000/3600 320∗10001∗3600 m/sec.. as 1 km =1000m and 1 hr = 3600 sec

OpenStudy (anonymous):

lets do it this way

OpenStudy (anonymous):

89.44 meters per second needs to convert to kilometers per hour

OpenStudy (anonymous):

so i do 1000/3600 320∗10001∗3600

OpenStudy (anonymous):

wait im confused

OpenStudy (anonymous):

so 89.44*60=5344.4*60=321984 that would give u meter per hour

OpenStudy (anonymous):

now to convert meters to kilometers

OpenStudy (anonymous):

it would be 321.984 kilometers per hour

OpenStudy (anonymous):

so for his train to be the second fastest it would have to be in between the two numbers 321.984 and 320 so round 321.984 to 322

OpenStudy (anonymous):

his train would have to go 321

OpenStudy (anonymous):

and to get the second part u plug in 321 kilometers per hour